# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v2

@@ 3.4 Two-dimensional Inviscid Burgers’ Equation
Figure 7 presents the spatial error plots in both training and extrapolation regimes, demonstrating that the time integrator models overall achieve higher accuracy than the baseline models. The solution plot for the same sample is presented in Figure A2. Figures 3(g) and (h) compare the relative $L_{2}$ error trends over the entire temporal domain. Notably, we reduced the training temporal domain to $t\in[0,0.33]$ with the extrapolation domain spanning $t\in[0.33,1.0]$, yielding the most challenging case in this work; all networks must perform $3\times$ extrapolation with no prior knowledge of shock formation at longer temporal horizons (no training data contained shocks for any framework). The autoregressive DeepONet rapidly accumulates error, exceeding 130% by the final time step. The full rollout DeepONet initially performs better, but its fixed-basis representation limits temporal generalization, with errors surpassing TI-DeepONet beyond $t=0.5$. Regarding the FNO variants, FNO FR fails catastrophically in capturing global system frequencies due to insufficient solution information across all timesteps, yielding completely unstable solution fields at longer temporal horizons with errors reaching 226% at the final timestep. FNO AR performs significantly better than DON AR, DON FR, and FNO FR, even outperforming the TI-based variants within the training temporal domain. However, due to its inherent sensitivity to error compounding during autoregressive inference, its excellent training domain accuracy (as low as 8% error) rapidly degrades in the extrapolation regime, reaching over 72% error at the final timestep, surpassing DON FR, which saturates at 52%. Consistent with previous examples, TI(L)-DeepONet maintains relatively slower and stable error growth, achieving final errors around 32% and outperforming all baselines as well as TI-DeepONet in the long-term prediction regime. These results demonstrate the robustness of time-integrator frameworks in modeling complex spatiotemporal stiff dynamics.
4 Choice of Inference Timestep
In our primary evaluations, TI-DeepONet was trained and tested using the same temporal resolution, i.e., the training and inference timesteps ($\Delta t$) were matched. However, the architecture’s compatibility with higher-order numerical integrators suggests that, in principle, inference may be conducted using timesteps different from those used during training. As long as the chosen $\Delta t$ satisfies the Courant–Friedrichs–Lewy (CFL) stability condition, the resulting time integration should remain stable. This flexibility stems from the fact that TI-DeepONet mirrors traditional numerical solvers: both employ explicit time-stepping schemes, with TI-DeepONet approximating the right-hand side of the PDE via a neural operator rather than finite-difference or finite-element discretizations of spatial derivatives.
To investigate this capability, we performed preliminary experiments on the 1D Burgers equation, evaluating TI-DeepONet at three inference timesteps, $\Delta t\in{0.1,0.01,0.001}$, while training was performed at $\Delta t=0.01$. Notably, TI-DeepONet preserved comparable $L_{2}$ error even when evaluated with a significantly coarser timestep ($\Delta t=0.1$) than that used during training. This finding suggests the potential for developing more efficient surrogates that accelerate inference without additional training costs.
5 Conclusions
Accurate long-term extrapolation of dynamical systems remains a fundamental challenge for neural operators, particularly due to error accumulation in autoregressive predictions. To address this, we introduce TI-DeepONet and its adaptive variant TI(L)-DeepONet, which integrates classical numerical time integration with an operator learning framework (DeepONet). By reformulating the learning objective to approximate instantaneous time derivatives-subsequently integrated via high-order numerical schemes, our frameworks preserve the causal structure of dynamical systems while enhancing stability and accuracy in long-term forecasting. Furthermore, the operator learning paradigm enables continuous spatiotemporal solutions for parametrized initial conditions, thereby enabling the construction of an efficient and reliable emulator for complex systems. Extensive evaluations across four canonical PDEs (1D/2D Burgers’, 1D KdV, and 1D KS equations) demonstrate the superiority of our approach. TI-DeepONet reduces relative $L_{2}$ extrapolation errors by 81% compared to autoregressive DeepONet and 70% compared to full-rollout methods, while maintaining stable predictions for temporal domains extending up to twice the training interval. The learnable coefficients in TI(L)-DeepONet further refine accuracy by adaptively weighting intermediate integration slopes, enabling solution-specific adjustments that outperform static numerical schemes in all our examples. By embedding temporal causality into the learning process and decoupling training-time integrators (e.g., Runge-Kutta) from inference-time integrators (e.g., Runge-Kutta or Adams-Bashforth), our methodology bridges neural operator flexibility with the robustness of numerical analysis. Among the proposed variants, TI(L)-DeepONet achieves the highest accuracy, albeit with a higher per-iteration computational cost due to its learnable integration coefficients (see SI Section A). However, this increase is offset by its significantly faster convergence to lower training loss, resulting in no net increase in total compute time. This hybrid paradigm offers a reliable pathway for modeling high-dimensional, time-dependent systems, where stability and generalizability are critical.
6 Limitations and Future Work
While our proposed frameworks demonstrate superior extrapolation accuracy compared to existing neural operator training strategies, they incur higher computational costs during both training and inference, as detailed in the SI Section A. Although these offline training costs can be justified by the consistent reliability of predictions across a wide range of parametric initial conditions, improving computational efficiency remains a key priority. To this end, we plan to leverage multi-GPU training and investigate methods for more efficient forward and backward propagation through the ODE solver. Additionally, we aim to extend the framework’s scalability to 3D and multiphysics systems and to incorporate adaptive time-stepping strategies for handling stiff dynamical systems.
Acknowledgments
Appendix A Limitations: Computational costs

### v3

@@ 3.4 Two-dimensional Burgers’ Equation
Figure 12 presents the spatial error plots in both training and extrapolation regimes, demonstrating that the time integrator models overall achieve higher accuracy than the baseline models. The solution plot for the same sample is presented in Figure A2. Figures 3(g) and (h) compare the relative $L_{2}$ error trends over the entire temporal domain. Notably, we reduced the training temporal domain to $t\in[0,0.33]$ with the extrapolation domain spanning $t\in[0.33,1.0]$, yielding one of the most challenging case in this work; all networks must perform $3\times$ extrapolation with no prior knowledge of shock formation at longer temporal horizons (no training data contained shocks for any framework). DON AR rapidly accumulates error, exceeding 130% by the final timestep. DON FR initially performs better, but its fixed-basis representation limits temporal generalization, with errors surpassing TI-DON beyond $t=0.5$. Regarding the FNO variants, FNO FR fails catastrophically in capturing global system frequencies due to insufficient solution information across all timesteps, yielding completely unstable solution fields at longer temporal horizons with errors reaching 226% at the final timestep. FNO AR performs significantly better than DON AR, DON FR, and FNO FR, even outperforming the TI-based variants within the training temporal domain. However, due to its inherent sensitivity to error compounding during autoregressive inference, its excellent training domain accuracy (as low as 8% error) rapidly degrades in the extrapolation regime, reaching over 72% error at the final timestep, surpassing DON FR, which saturates at 52%. Consistent with previous examples, TI(L)-DON maintains relatively slower and stable error growth, achieving final errors around 32% and outperforming all baselines as well as marginally outperforming TI-DON in the long-term prediction regime. A more detailed investigation is presented through the test error distribution plots in Figure 13, where the overall error distributions for TI-DON and TI(L)-DON show comparable performance on a per-sample basis. Among the baseline methods, FNO AR demonstrates largely comparable performance to the TI-variants, followed by DON FR, while FNO FR and DON AR exhibit similar but less favorable error distributions. These results demonstrate the robustness of time-integrator frameworks in modeling complex spatiotemporal dynamics involving stiff gradients and shock formation. Finally, we present histograms of the optimally learned weighting coefficients $\alpha_{i}$ for $i\in\{1,2,3,4\}$ in the TI(L)-DON framework and compare them against the classical RK4 weighting coefficients in Figure 14. Interestingly, $\alpha_{3}$ exhibits values close to the classical RK4 coefficient of 0.33, while $\alpha_{4}$ and $\alpha_{2}$ lie within 0.04 and 0.1 respectively, indicating that they are less significant compared to the other coefficients for this problem. In contrast, $\alpha_{1}$ shows values greater than 0.5, substantially higher than the classical coefficient of 0.17. Overall, for the 2D Burgers’ case, we observe that $\alpha_{3}$ and $\alpha_{1}$ are the dominant coefficients, with the latter bearing the most significance. These observations reinforce a key insight: the learned weighting coefficients are data-dependent and tailored to the underlying solution dynamics provided as input to the auxiliary RK4 network. This example reveals that for problems where shock formation is imminent, the use of such adaptive schemes is particularly advantageous, as they provide instantaneous solution-specific adaptivity that fixed-coefficient schemes cannot offer.
Figure 15: 2D Burgers’ Equation: Timestep refinement study with different inference $\Delta t$ for the TI-based variants.
As described in Fig. 2, the architecture of TI-DON allows for flexibility in the choice of different $\Delta t$ at training and inference. To this end, we perform a timestep refinement study for the 2D Burgers’ equation, where the model was trained with $\Delta t=0.01$ and inference was performed using $\Delta t\in\{0.1,0.05,0.01,0.005\}$. The relative $L_{2}$ error growth is presented in Fig. 15 for both TI-DON and TI(L)-DON with different inference $\Delta t$. For TI-DON, prediction errors remain consistent across all $\Delta t$ values, demonstrating timestep-invariant behavior and confirming that error growth stems from model approximation rather than numerical instability. For TI(L)-DON, finer temporal resolutions ($\Delta t\in\{0.01,0.005\}$) yield marginally reduced errors at extended horizons, indicating that this method potentially benefits from smaller timesteps.
3.5 Two-dimensional Scalar Rotation Advection-Diffusion
The fifth example that we consider is the scalar advection–diffusion equation with a rotating advection field as investigated in de Castro et al. (2023). The governing PDE reads as:
$\dot{u}_{i}-\nabla\cdot F_{i}(u_{i})$ $=f_{i}$ $\text{on }\Omega_{i}\times[0,T],$ (11)
$u_{i}$ $=g_{i}$ $\text{on }d\Omega_{i}\times[0,T]~~i=1,2,$
$u_{i}$ $=u_{i,0}$ $\text{on }\Omega_{i},~t=0$
where $\dot{u_{i}}$ refers to the instantaneous time-derivative of the scalar field $u(\mathbf{x},t)$ at the $i^{th}$ state, $F_{i}(u_{i})=\kappa_{i}\nabla u_{i}-\mathbf{a}u_{i}$ is the total flux function, $f_{i}=f_{i}(\mathbf{x},t)$ is a source term, $g_{i}=g_{i}(\mathbf{x},t)$ is the prescribed boundary condition, $u_{i,0}=u(\mathbf{x},t=0)$ is the defined initial condition, $\kappa_{i}=\kappa_{i}(\mathbf{x},t)>0$ is the diffusion coefficient in the spatial domain $\Omega_{i}$, and $\mathbf{a}=\mathbf{a}(\mathbf{x},t)$ is the rotating advection field.
Data generation and preparation:
For this study, we consider the computational domain $\Omega=(0,1)^{2}$ discretized using a uniform $64\times 64$ quadrilateral mesh. The rotating advection field is defined as $\mathbf{a}=(0.5-y,x-0.5)$, which induces rigid body rotation about the domain center. A constant diffusion coefficient $\kappa=0.001$ is employed, with homogeneous Dirichlet boundary conditions prescribed on all boundaries. The initial conditions consist of randomly parameterized Gaussian blobs with amplitude $A$, center coordinates $(x_{0},y_{0})$, and standard deviations $\sigma_{1},\sigma_{2}$, all uniformly sampled from the intervals $A\in[0.3,1.0]$, $(x_{0},y_{0})\in[0.2,0.8]^{2}$, and $\sigma_{1},\sigma_{2}\in[0.03,0.10]$, respectively. The governing equations are solved using a streamline upwind Petrov-Galerkin stabilized finite element method (SUPG) with Crank-Nicolson time integration. The dataset comprises 800 samples, each simulated over a quarter rotation period ($T=\pi/2$), with solution snapshots saved at regular intervals of $\Delta t\approx 0.0067$, resulting in a total of 101 timesteps. For further details regarding the problem formulation, the reader is referred to de Castro et al. (2023). The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=640$ training samples and $N_{\text{test}}=160$ testing samples. The dataset arrangement for full rollout, autoregressive, and TI-based methods follows the same protocol as described in Section 3.4, with the training temporal domain $t_{\text{train}}\in[0,0.27]$ (first 40 out of 101 timesteps) and extrapolation domain $t\in[0,0.67]$.
Training architecture:
Tables LABEL:tab:2d_rotating_advection_diffusion_don-architecture and A11 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size and Fourier modes (if applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the DON FR, the spatial and temporal coordinates $(x,y,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 3 to 63. For the autoregressive and TI-based variants, only the spatial coordinates $(x,y)$ are augmented with Fourier features $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky)]$, expanding the input dimension from 2 to 42. In this case, for the TI(L)-DeepONet architecture, we employ a similar combination of 2D convolutional blocks and MLP layers as used in the branch network to learn the RK4 slope coefficients. The convolutional blocks utilize the same combination of Conv2D and pooling layers as the branch, with the only difference being that the Conv2D layers have a reduced number of filters (32 instead of 64). This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers and finally to an output layer with 4 neurons that predicts the RK4 slope coefficients. The Conv2D layers use a GELU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted slope coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 steps. The auxiliary network responsible for predicting the learnable RK4 slope coefficients was trained with a similar exponential decay schedule, but with a higher initial learning rate of $4\times 10^{-3}$. For both FNO AR and FNO FR variants, we employ a similar exponential decay schedule as the DeepONet variants albeit with a decay factor of 0.96 every 2000 steps.
Figure 16: 2D Scalar Rotation Advection-Diffusion: Spatial error distribution across training ($t\in[0,0.27]$) and extrapolation ($t\in[0.27,0.67]$) regimes for all frameworks, illustrated with a representative sample.
Figure 17: 2D Scalar Rotation Advection-Diffusion: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
Following a similar analysis as the previous cases, a qualitative comparison of the performance of the various frameworks is presented through the spatial error distribution plots in the interpolation ($t\in[0,0.27]$) and extrapolation ($t\in[0.27,0.67]$) regimes in Figure 16. DON AR incurs the highest errors due to its greater sensitivity to error accumulation in both the interpolation and extrapolation domains, thereby starting to accumulate errors early in the temporal evolution. This is followed by FNO FR, which, consistent with previous cases, suffers from incomplete spectral representation due to computing Fourier features over a limited temporal domain, and therefore incurs errors in both the interpolation and, more prominently, the extrapolation regime. Next is DON FR, which exhibits impressive accuracy during interpolation since its learned basis is tailored to the training temporal domain, but immediately breaks down upon entering the extrapolation regime. The TI-based variants follow, exhibiting much more stable error growth and demonstrating excellent performance not only in interpolation but also in extrapolation, owing to the physics prior induced by numerical time integration. However, among all baselines, FNO AR emerges as the most dominant method, exhibiting errors an order of magnitude lower than other baselines ($4\%$), while the second-best error is approximately $23.72\%$ achieved by TI(L)-DON at the final timestep. The results presented in Table 1 and Figure 3 are in strong agreement with these qualitative trends. The error growth of DON AR is, as expected, monotonic due to compounding of model approximation errors during recursive inference and surpasses FNO FR at $t=0.13$. FNO FR incurs a non-monotonic yet increasing error growth throughout the temporal domain, following DON AR, with both methods achieving $L_{2}$ errors above 80-97% at the final timestep. DON FR performs well in interpolation as highlighted above and achieves comparable performance to TI-DON for $t\in[0.13,0.27]$, but rapidly degrades upon entering the extrapolation domain, ultimately reaching errors over 105% and surpassing all other baselines in performance degradation. Compared to DON FR, DON AR, and FNO FR, TI-DON’s inherent stabilization of error growth leads to roughly 23.7% error at the final timestep, which is eventually dominated by the impressive performance of FNO AR, which exhibits a plateau in error growth during the later timesteps in the extrapolation domain. As evidenced in Figure 3(j) and Figure 17, there is an insignificant difference in the error growth as well as overall error distributions between TI-DON and TI(L)-DON, with both exhibiting similar performance. Therefore, for brevity, we do not analyze the distribution of the learned $\boldsymbol{\alpha}$ coefficients for the intermediate RK4 slopes in the TI(L)-DON framework for this case. Moreover, the evidently superior performance of FNO AR over all the other NO baselines is solidified, with errors one order of magnitude less than the TI-based variants. The rest of the ranking is as follows: FNO FR $\approx$ DON FR >> DON AR. This example elucidates two key conclusions. First, the rotation advection-diffusion problem poses a more challenging test case compared to the previous examples. Given the negligible diffusion coefficient ($\kappa=0.001$), the problem is strongly advection-dominated, and the Gaussian blob should undergo near-rigid body rotation with its shape preserved throughout the temporal evolution. However, the TI-based variants incur noticeable errors in capturing this rotational dynamics. Specifically, the sequential predictions distort the boundaries of the Gaussian blob, violating the shape-preserving constraint inherent to rigid body rotation. This artifact accumulates over successive time integrations, leading to progressive deformation of the blob structure. We attribute this to the sensitivity of explicit time integration schemes to the choice of $\Delta t$ when resolving rotational motion; a finer temporal resolution may be necessary to minimize boundary distortions and achieve performance comparable to FNO AR. The superior performance of FNO AR in this case can be attributed to the global Fourier basis, which naturally captures the periodic structure inherent in rotational dynamics without requiring explicit temporal marching, thereby avoiding the accumulation of shape-distorting artifacts. Second, this example highlights that for certain dynamical systems, the choice of neural operator architecture may be as important as the time integration strategy. This motivates a more general TI-NO formulation (time-integrator embedded neural operator), wherein the neural operator and time integration scheme can be selected independently based on the problem characteristics. We hypothesize that a TI-FNO variant, which combines the time-integrator framework with FNO’s spectral representation, could leverage the strengths of both approaches for problems involving rotational or periodic dynamics. Such an extension remains beyond the scope of the present work and is deferred to future investigation.
3.6 Three-dimensional Heat Conduction
We consider three-dimensional heat conduction as the final example in this study. The heat equation is a linear parabolic PDE with a wide range of applications: it naturally arises in modeling heat transfer through a medium, appears in probability theory as the governing equation for Brownian motion, plays a fundamental role in mathematical finance through option pricing models, and is widely used in image processing for smoothing and denoising. In a three-dimensional domain, the governing equation is given by:
$\frac{\partial T}{\partial t}=\alpha\left(\frac{\partial^{2}T}{\partial x^{2}}+\frac{\partial^{2}T}{\partial y^{2}}+\frac{\partial^{2}T}{\partial z^{2}}\right),\quad\forall\ (x,y,z,t)\in[0,1]^{3}\times[0,1],$ (12)
where $\alpha=1.0$ denotes the thermal diffusivity. Here, $T$ represents the scalar temperature field with initial condition $T_{0}$.
Data generation and preparation:
The spatial domain is a three-dimensional L-shaped region constructed by removing the upper-right quadrant (where $x\geq L_{x}/2$ and $y\geq L_{y}/2$) from a rectangular block, with the L-shaped cross-section oriented in the XY-plane. The 3D heat conduction equation is solved using an explicit FDM solver on a $32\times 32\times 16$ Cartesian grid over the spatial domain $(x,y,z)\in[0,1]^{3}$. The Laplacian term on the RHS is discretized using a second-order central difference (CD2) scheme. Based on von Neumann stability analysis, a timestep of $\Delta t=0.01$ is chosen, and the solution is advanced in time up to $t=1.0$. The initial condition consists of a Gaussian blob with standard deviation $\sigma=4.0$ grid points. The amplitude of the Gaussian is uniformly sampled from the interval $[0,1]$. The center of the Gaussian $(x_{c},y_{c},z_{c})$ is uniformly sampled within the lower-left block of the L-shaped domain, with a margin of $\sigma$ from all boundaries to ensure the blob remains fully contained. Specifically, the center coordinates are sampled as $(x_{c},y_{c},z_{c})\in[4,11]^{3}$ in grid indices. Homogeneous Dirichlet boundary conditions ($T=0$) are imposed on all faces of the L-shaped block. A total of 1000 samples are generated with an 80-20 train-test split. The training temporal domain spans $t\in[0,0.33]$, while the extrapolation domain covers $t\in[0.33,1.0]$. The data preparation strategy follows the same protocol as the previous PDE examples for all neural operator baselines. For further details regarding the data generation procedure, the reader is referred to Roy et al. (2025), from which the data generation code is adapted.
Training architecture
Table LABEL:tab:3d_heat_don-architecture presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (if applicable). Unlike previous cases, the trunk network is not equipped with Fourier feature encodings, as sufficiently accurate convergence was observed without them. For the TI(L)-DON architecture, the auxiliary network for predicting the RK4 weighting coefficients employs Conv3D layers with 32 filters, analogous to those in the branch network. This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers and finally to an output layer with 4 neurons that predicts the RK4 weighting coefficients. The Conv3D layers use a GELU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.96 every 2000 steps. The auxiliary network was trained with a similar exponential decay schedule but with a higher initial learning rate of $5\times 10^{-3}$.
Figure 18: 3D Heat Conduction: Schematic showing the XY, YZ, and XZ slicing planes (from left to right) for visualizing the error and solution profiles.
Figure 19: 3D Heat Conduction: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated using a representative sample. The contours are presented along the XY slicing plane (see Fig. 18 for details regarding the location of the XY slice).
Figure 20: 3D Heat Conduction: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated using a representative sample. The contours are presented along the YZ slicing plane (see Fig. 18 for details regarding the location of the YZ slice).
Figure 21: 3D Heat Conduction: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated using a representative sample. The contours are presented along the XZ slicing plane (see Fig. 18 for details regarding the location of the XZ slice).
Figure 22: 3D Heat Conduction: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
Figures 19, 20, and 21 present the error contours obtained for the DeepONet-based variants, i.e., TI(L)-DON, TI-DON, DON FR, and DON AR, in the interpolation ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1.0]$) regimes for a representative test sample, visualized along different slicing planes: XY, YZ, and XZ (in that order). FNO-based baselines are not shown here, as they cannot be directly applied to irregular domains such as the L-shaped domain considered in this study. As expected, DON AR exhibits an order of magnitude higher error due to its increased sensitivity to error accumulation. During sequential predictions, the compounding of model approximation errors causes DON AR to “artificially” dissipate the Gaussian blob too rapidly; consequently, the blob is visibly absent from the initial few timesteps itself (see Figures A4, A5, and A6). This is superseded by the performance of DON FR, which demonstrates high accuracy in the interpolation regime but quickly deteriorates in the extrapolation regime due to its limited fixed-basis representation capacity tailored to the training temporal domain. In contrast, the TI-based variants consistently achieve superior performance in both the interpolation and extrapolation regimes by mitigating error accumulation through the physics-aware numerical time integration module. These qualitative trends are in good agreement with the quantitative results of relative $L_{2}$ error growth during inference, as shown in Table 1 and Figure 3. Notably, compared to DON AR and DON FR, which incur relative $L_{2}$ errors as high as 0.98 and 0.22 at the final timestep respectively, the reduction in error achieved by the TI-based variants is remarkable; approximately 96.3% and 83.6% improvement with respect to their autoregressive and full rollout counterparts. This example serves as a benchmark demonstrating that, especially in high-dimensional spatiotemporal dynamics (3D in this case), where spatial errors along three dimensions can compound uncontrollably, TI-based variants provide a pathway for mitigating error growth and stabilizing extrapolation owing to the strong physics-aware inductive bias imposed by the numerical time integration module. Unlike autoregressive DON, the diffusive characteristics of the heat conduction equation are extremely well captured by the TI-based variants, with TI(L)-DON slightly outperforming TI-DON, as evidenced in Figure 22. This is followed by the performance of DON FR and then DON AR.
4 Computational Costs